• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name minakat

    Log entry time 23:33:31 on December19,2007

    Entry number 222305

    keyword=Left IOC/run plan update

    Tech is in the hall trying to fix left IOC problem.
    As per J.P.'s instructions we are to scratch the remainder
    of the boiling tests and proceed with the optics tests.

    Setting the right spectrometer to first optics point.

    Set the optics 7 foil target in.

    If IOC problem persists (again as per J.P.'s instructions)
    we will ask the tech to leave the hall and run only the
    right spectrometer and be in production by
    00:30 at the latest.